(3x+1)dx / ((x+1)^12(3x-1)^12)


1
Expert's answer
2021-02-24T07:37:30-0500

Solution:

Let "I=\int \dfrac{3x+1}{(x+1)^{12}(3x-1)^{12}}dx"

"=\int \dfrac{3x+1}{[(x+1)(3x-1)]^{12}}dx"

"=\int \dfrac{3x+1}{(3x^2+2x-1)^{12}}dx" ...(i)

Now put "(3x^2+2x-1)=t" ...(ii)

On differentiating w.r.t. "x"

"6x+2=\dfrac{dt}{dx}"

"\Rightarrow (3x+1)dx=\dfrac{dt}{2}" ...(iii)

Putting (ii) and (iii) in (i), we get

"I=\int \dfrac{1}{(t)^{12}}\dfrac{dt}{2}"

"=\dfrac{1}{2}\int \dfrac{1}{t^{12}}dt"

"=\dfrac{1}{2}\int {t^{-12}}dt"

"=\dfrac{1}{2}(\dfrac{t^{-12+1}}{{-12+1}})+C"

"=\dfrac{1}{2}(\dfrac{t^{-11}}{{-11}})+C"

"=\dfrac{-1}{{22t^{11}}}+C"

"=\dfrac{-1}{{22(3x^2+2x-1)^{11}}}+C" [Using (ii)]

Thus, our final answer is "\dfrac{-1}{{22(3x^2+2x-1)^{11}}}+C"